### v3

# Modeling Issues with Eye Tracking Data

Abstract
I describe and compare procedures for binary eye-tracking (ET) data. These procedures are applied to both raw and compressed data. The basic GLMM model is a logistic mixed model combined with random effects for persons and items. Additional models address autocorrelation eye-tracking serial observations. In particular, two novel approaches are illustrated that address serial without the use of an observed lag-1 predictor: a first-order autoregressive model obtained with generalized estimating equations, and a recurrent two-state survival model. Altogether, the results of five different analyses point to unresolved issues in the analysis of eye-tracking data and new directions for analytic development.
keywords:
eye-tracking modeling generalized linear mixed model data compression generalized estimating equations survival analysis autocorrelation
1 Introduction
Modeling the cognitive processes with which students engage in psychological tasks in eye tracking studies has important implications for theory development. It may also enable the construction of more informative test scores [DeBoeck2019]. A number of models have been used for the analysis of eye-tracking data [Barr2008, Chen2019, BrownSchmidt2020]. The basic features of these models include binary or categorical data for eye fixations, and multiple measurements of subject responses over time. The data then manifest as subject-level time series, often composed of Bernoulli observations. [Cho2018] extended the original work of [Barr2008] using a generalized linear mixed model (GLMM) to control autocorrelation, a common feature of time series that may bias standard errors. In this paper, I also present an alternative procedure that features run length encoding (RLE) combined with a two-state survival model.
After a brief review of previous studies, two methods of compressing data for eye-tracking analyses are described. In part, compression leads to more efficient analysis and can reduce observed serial correlation in eye-tracking time series. Then I describe six methods for analyzing eye-tracking data, where a method refers to a statistical model combined with a set of raw or compressed data. After analyses are completed, model coefficients are reported with primary attention placed on the experimental effects. Discussion follows concerning further directions for modeling in educational eye-tracking studies.
2 Relevant Literature
The original multilevel model for analyzing eye-tracking data was constructed by [Barr2008] in terms of a cross-classified (subjects by items) design. This logistic regression model included fixed effects for experimental variables. [Ryskin2015] applied a mixed-effects logistic regression model to cross-classified eye-tracking data to examine how listeners use a speaker’s perspective to identify referents during unscripted conversational instructions. Several studies have updated mixed effects modeling to include effects for autocorrelated data and time ([cho2018ar, BrownSchmidt2020, ChoTree, BROWNSCHMIDT2025149511, Cho_Brown-Schmidt_Clough_Duff_2025]).
The subject observations analyzed in the current study are eye-tracking fixations for visual targets, originally collected by [Ryskin2015]. The hierarchical structure of the data is characterized by trial-level data nested within subjects and items. For a visual presentation of the data structure, see Figure 3 of [Cho2018]. On each trial, the speaker and the listener participated together in a brief conversational exchange aimed at identifying a target object (the different objects constitute a set of items). Although they were seated in separate rooms, they communicated naturally through wireless microphones while their eye movements were recorded simultaneously with two remote eye trackers. The variable $Contrast$ captures whether the speaker uniquely identified the target object or whether multiple objects in the listener’s visual scene could match the speaker’s descriptions. It reflects the amount of referential competition the listener must cognitively manage during real-time interpretation. Referential competition was hypothesized to make the task more difficult. The variable $Privileged$ reflects whether listeners could reduce referential competition based on information they infer from the speaker’s perspective; this variable was hypothesized to make the task easier.
[Cho_Brown-Schmidt_Clough_Duff_2025] addressed several issues with GLMM models in eye-tracking studies related to how observations are structured. In particular, binary time series are generated as runs of 0s and 1s, and longer gazes generate longer runs as do higher sampling rates. However, longer runs necessarily result in higher autocorrelation which can affect the quality of estimates and standard errors. For this reason, [Cho_Brown-Schmidt_Clough_Duff_2025] constructed a generalized mixed model (GLMM) with a lag-1 outcome variable as a predictor in addition to other covariates. This resulted in estimating an autoregression coefficient AR(1) for observed data. Random effects were also included for subject and items as well as the AR(1) coefficient to account for differing levels of autocorrelation by subject and by item. [Cho_Brown-Schmidt_Clough_Duff_2025] showed that autoregressive GLMM models can recover parameters accurately in a simulation study, but they noted that "alternative methods for modeling AR effects in binary time series data are still in the nascent stages" (p. 652). The goal of this paper is to contribute to this effort.
3 Data
The data structure is organized around repeated measures, where eye-tracking observations are collected for subject–item–trial combinations. Each subject completed multiple trials, each associated with a visual display [Ryskin2015]. Each trial included two perspective-taking conditions ($Contrast$ and $Privileged$) as previously described; these two conditions were assigned as the item level and thus varied across items. In the original data, observations consisted of $n=112$ observations at .01 second intervals. Data are recorded as the number of discrete instances in which a person is gazing at a target area (1) versus off-target (0). The total number of data points usually manifests as strings of 1s and 0s. The original data were presumably recorded at a frequency of 1000Hz and compressed to .01 second intervals. The resulting data file consisted of 995,232 records in long format for 152 subjects.
3.1 Run Length Encoding
Consider an example string of 10 samples of observations taken at equal intervals: 1110001111, where each sample fixation is .01 second long. (Note this is a simplification because recorded data depend to some degree on how data are compressed by the eye-tracking apparatus prior to data file construction.) Assume recorded instances of 1 (on target) and 0 (off target) are collected beginning at start time 0 and ending at 1 second, so that each 0-1 value corresponds to a .1 second time lapse.
Table 1: Run Length Encoding
Run y n StopTime
1 1 3 0.03 sec
2 0 3 0.06 sec
3 1 4 0.10 sec
run length encoding (RLE) is illustrated in Table 1. The 10 records are compressed into three, which would seem like a loss of information. However, knowing the bin size is .01 seconds, or the total time interval and number of bins, would allow the original data set to be reconstructed in its entirety. This suggests there is no more or less information in the compressed than in the original data. Thus, RLE is lossless compression. The RLE data are used with the survival model described below.
3.2 Data Selection for Analysis
I modified the data by excluding two types of subject response. First, in the original study there were multiple trials in which the same item could be presented more than once for a subject. I retained only the first trial of an item for each subject. Second, there were subject-by-item time series with only a single gaze recorded. In many of these instances, the subject appeared to be uniformly gazing off-target for two 2-3 items in a row. This type of data is commonly considered "missing" in eye-tracking studies, which is related to long off-screen periods [rayner2009eyemovements, staub2013individual]. These particular series were deleted from the data in the current study. This selection procedure resulted in an uncompressed long-format data set of 646,016 records. An RLE data set was then constructed as a long-format data set of 15,025 records (a compression rate of about 98%).
4 Modeling
Modeling procedures in this study are applied to the data collected by [Ryskin2015] and previously analyzed by Cho et al. (2018). The data described above were analyzed by [Cho2018] using a generalized linear mixed model. This model included a lag-1 observed outcome as a predictor to account for autocorrelation. Other predictors included experimental variables. The model also incorporated random effects for subjects and items. In this paper, I use several similar models for data obtained with the exclusions described previously. I also added the two interaction terms TimePrivilegedTime*Privileged and TimetTime*t. These interactions are designed to detect whether the effects of the experimental variables change across time.
The first model was run on the full uncompressed data set. I refer to this as the GLM model, which is given by:
$Y_{ijt}$ $\sim\text{Bernoulli}\!\left(p_{ijt}\right),$ (1)
$\text{logit}\!\left(Y_{ijt}=1|p_{ijt}\right)$ $=\mu+\beta_{1}\,\text{Contrast}_{ij}+\beta_{2}\,\text{Privileged}_{ij}$
$\quad+\beta_{3}\,\text{Time}_{t}*\text{Contrast}_{ij}+\beta_{4}\,\text{Time}_{t}*\text{Privileged}_{ij}$
$\quad+u_{i}+v_{j},\text{where}$
$u_{i}$ $\sim N(0,\sigma_{u}^{2}),$
$v_{j}$ $\sim N(0,\sigma_{v}^{2}).$
Here, the binary response $y_{ijt}$ for subject $i$ and item $j$ at time $t$ is modeled using a Bernoulli regression with fixed effects for t, Privileged, the interaction of $Time$ with $Contrast$ and $Privileged$, and random effects for subject and item. The LAG model was obtained by augmenting the GLM model with a lag-1 predictor $y_{ij,t-1}$ to account for serial recursion:
$Y_{ijt}$ $\sim\text{Bernoulli}\!\left(p_{ijt}\right),$ (2)
$\text{logit}\!\left(Y_{ijt}=1|p_{ijt}\right)$ $=\mu+\beta_{0}\,\text{y}_{ij,t-1}+\beta_{1}\,\text{Contrast}_{ij}+\beta_{2}\,\text{Privileged}_{ij}$
$\quad+\beta_{3}\,\text{Time}_{t}\!*\!\text{Contrast}_{ij}+\beta_{4}\,\text{Time}_{t}\!*\!\text{Privileged}_{ij}$
$\quad+u_{i}+v_{j},$
where the random effects $u$ and $v$ share the same variance structure in Equation 2.
The GLM model is altered more substantially by omitting random effects and incorporating a coefficient $\phi_{A}$ as an alternative for the lag-1 autocorrelation. Error correlation is then incorporated directly into the variance structure. To show the application of this approach, generalized estimating equations can be used. To show the application of this structure, it is expanded in some detail.
5 Generalized Estimating Equations.
Generalized Estimating Equations (GEE) [LiangZeger1986] provide a semiparametric approach for marginal regression with correlated errors. For binary outcomes, let $p_{ijt}=\Pr(Y_{ijt}=1)$ denote the marginal mean for subject $i$ for item $j$ at time $t$. The mean model is written as
$Y_{ijt}$ $\sim\text{Bernoulli}\!\left(p_{ijt}\right),$ (3)
$\text{logit}\!\left(Y_{ijt}=1|p_{ijt}\right)$ $=\mu+\beta_{1}\,\text{Contrast}_{ij}+\beta_{2}\,\text{Privileged}_{ij}$
$\quad+\beta_{3}\,\text{Time}_{t}\ *\text{Contrast}_{ij}+\beta_{4}\,\text{Time}_{t}\ *\text{Privileged}_{ij}$
The marginal variance of each Bernoulli observation is $\sigma_{ijt}^{2}=\sigma^{2}(Y_{ijt})=p_{ijt}(1-p_{ijt})$, and the variance-function matrix for cluster $ij$ is $A_{ij}=\mathrm{diag}(\sigma_{ij1}^{2},\ldots,\sigma_{ijT}^{2})$, where $T=112$. Serial or within-cluster correlation is introduced by specifying a working correlation matrix $R$ containing the single parameter $\phi_{A}$. For AR(1) structure, $R$ is a symmetric matrix with off-diagonal elements defined as $\phi_{A}^{|l-m|}$, where $l$ and $m$ are two time indices. The observed vector $Y_{i}=(Y_{ij1},\ldots,Y_{ijT})^{\top}$ then has the variance structure:
$V_{ij}$ $=\alpha A_{ij}^{1/2}\,R\,A_{ij}^{1/2},$ (4)
where $\alpha$ is the scale parameter. Thus, the residual correlation $\phi$ enters the estimating equations through $R$.
Let
$D_{ij}=\frac{\partial p_{ij}}{\partial\beta^{\top}}$
be the $T\times q$ matrix of derivatives of the mean vector (where $q=6$) with respect to $\beta$. The GEE estimator $\hat{\beta}$ is then obtained by solving the quasi-score equation
$\sum_{i=1}D_{ij}^{\top}V_{ij}^{-1}(Y_{ij}-p_{ij})=0.$ (5)
Estimation of the working correlation parameter $\phi$ is performed by a moment equation based on Pearson residuals,
$r_{it}=\frac{Y_{ijt}-p_{ijt}}{\sigma_{ijt}},\qquad r_{ij}=(r_{ij1},\ldots,r_{ijT})^{\top}.$
For an $\mathrm{AR}(1)$ structure, a common estimator is
$\hat{\phi}=\frac{\displaystyle\sum_{ij}\sum_{t=2}^{T}r_{ijt}\,r_{{ij},t-1}}{\displaystyle\sum_{ij}(T-1)},$ (6)
where $\hat{\phi_{A}}$ is the residual correlation, and the large-sample covariance of $\widehat{\beta}$ is estimated using the robust sandwich estimator is
[…]
$B$ $=\sum_{ij}D_{ij}^{\top}V_{ij}^{-1}D_{ij},$
$M$ $=\sum_{ij}D_{ij}^{\top}V_{ij}^{-1}(Y_{ij}-\hat{p}_{ij})(Y_{ij}-\hat{p}_{ij})^{\top}V_{ij}^{-1}D_{ij}.$
In Equations (5) and (7) it can be seen that $\phi_{A}$ affects both fixed-effect estimates and standard errors.
Several authors have noted that the GEE working covariance $V_{i}=A_{i}^{1/2}R_{\alpha}A_{i}^{1/2}$ plays a role analogous to the autocorrelation of latent residual processes in GLM or state-space models for binary data. In particular, [ZegerLiangAlbert1988] demonstrated that serial correlation among binary observations can be represented either by specifying a latent Gaussian process with AR(1) correlation or by imposing the same AR(1) structure on the working correlation matrix in GEE. Related comparisons appear in [HeagertyZeger2000], who show that marginal models in GEE and latent-process conditional models often share identical first- and second-moment structures, though they differ in higher-order behavior and interpretation.
Rather than controlling for autocorrelation with an observed lag-1 indicator, $\phi_{A}$ can me incorporated directly into the model’s variance structure. Residual structures are more likely to represent a stationary process, especially when $Time$ is included as a covariate. [Levin2018ChangeBlindness] previously applied a GEE model to analyze eye-tracking data; however, they used an unstructured correlation matrix as opposed to a first-order error structure. The AR1 variance structure achieves a higher degree of efficiency by estimating a single coefficient $\phi_{A}$, but this economy depends on the length of the time series.
A related model to AR1 is a first-order moving average model MA1. When binary data is composed of runs (streaks of 0s and 1s), it suggests that once a saccade occurs (e.g., a state switch), its influence is immediate but might not persist indefinitely across the entire time series. This seems more natural to a eye-tracking time series than an autogressive process. This model can also be fit with GEE in the forms of an $R$ matrix having a tri-diagonal structure, also known as a banded Toeplitz matrix of width 1. For this correlation matrix, only the off-diagonals immediately above and below the main diagonal have a nonzero value of $\phi_{M}$. The estimating equations are identical to those of the AR1 model.
6 Survival Modeling
RLE modeling of eye tracking responses is designed to model the 0-1 category frequencies, rather than the 0-1 data individually. This strategy was suggested previously in the discussion of Table 1. For this purpose, binomial modeling can be used in the framework of a GLMM. However, it is convenient to implement this analysis relative to runs within a subject-item cluster. In fact, the runs are treated as discrete instances of gaze that do not correspond to a linear process, either manifest or latent.
For RLE compression, the blocks correspond to $Y=0$ or to $Y=1$. After compression, each episode can be represented by an interval $Time_{\text{start}},Time_{\text{stop}}]$ and an associated event indicator that marks whether a transition to the opposite gaze state occurred at the end of the interval. This representation lends itself to an alternating two–state survival framework, where separate transition hazards are modeled for leaving the target $1\rightarrow 0$ and entering the target $0\rightarrow 1$ [kleinbaum2012survival, therneau2024survival]. Because RLE compression preserves the dwell structure, the resulting episode durations provide suitable survival intervals for Cox regression. The COX model is:
$h(t\mid X,S)$ $=h_{0,S}(t)\,\exp\Bigl(\beta_{1}\,\text{Privileged}+\beta_{2}\,\text{Contrast}$ (8)
$\quad+\beta_{4}\,\{0\!\to\!1\}*\text{Privileged}$
$\quad+\beta_{5}\,\{0\!\to\!1\}*\text{Contrast}\Bigr)$
In a recurrent two–state process with states $\{0,1\}$, the hazard function $h_{S\to S^{\prime}}(t)$ describes the instantaneous rate of transition from state $S$ to state $S^{\prime}$ at time $t$, conditional on the process having remained in state $S$ up to time $t$. Equivalently, the hazard quantifies how quickly a transition is expected to occur at time $t$, given that the system is currently in state $S$.
In the fitted model, the baseline hazard is stratified by transition type so that the temporal risk profiles for leaving and entering the target object may differ freely. Predictors such as Privileged and t are allowed to influence each transition differently through interaction terms with the transition indicator. This structure has interpretable effects: the main coefficients describe how covariates affect the hazard of leaving the target, while the interaction coefficients quantify how those effects change during target reacquisition. Main and transition effects and standard errors can be combined to represent total effects. For this analysis, the main focus is on the total effect for the $0\rightarrow 1$ transition. The survival model with RLE data provides a computationally efficient modeling alternative to AR1 with uncompressed data.
The model incorporates a Cox proportional hazards framework, the process being modeled is not survival in the conventional sense. The data consist of a binary time series in which each subject–item trajectory alternates repeatedly between two states, with each interval representing the duration spent in a given state before switching to the other. In this context, the COX model is used as a flexible estimator of state-specific switching intensities rather than time to a terminal event. By stratifying on transition direction, separate baseline hazards arfe used for $0\rightarrow 1$ and $1\rightarrow 0$ transitions, so that covariate effects are interpreted as changes in the instantaneous rate of leaving the current state.
7 Results
Analyses of the GLM and LAG methods were carried out with the R function glmer from the package lme4 [Bates2015_lme4]. The R function glmgee from the R package glmgee [Vanegas2023] was used for analysis of the AR1/MA1 models, and the R function coxph from the package survival was used for the two-state COX model [therneau2024survival]. The AR1 and MA1 models produced positive definite $R$ matrices; however, their determinants were near zero. To address this issue, user-defined $R$ matrices were constructed with $\phi=.95$ and a ridge of 1e-5 was applied to $R$. In both cases, estimates were nearly identical to those produced in runs where $\phi$ was freely estimated.
Because of the data exclusion procedures used in the present study, analytic results cannot be directly compared to prior studies. In this section, only the models are evaluated. The data for the models was modified in two ways. First, models using a lag-1 predictor require dropping the first observation in each individual time series. Second, for the survival model, the last run was deleted. Because the model focuses on transitions, the transition indicator was missing by fiat for the last run. To account for non-independence across subject and items, and a subjects-by-item indicator was used as a cluster variable in the AR1, MA1, and COX models.
Table 2: Fixed-effect estimates for five methods.
Term GLM AR1 MA1 LAG COX
Intercept -0.437 -0.393 -0.400 -2.946
Ylag-1 5.784
Privileged 0.077 0.052 0.070 0.083 0.053
Contrast -0.079 -0.057 -0.068 -0.013 -0.046
Time 0.854 0.813 0.831 0.498
Priv*Time -0.012 0.001 0.006 0.012
Contrast*Time -0.015 -0.005 -0.012 0.028
  • Note.Total effects are reported for the COM model.
7.1 Fixed Effects
The estimates of fixed effects for these models are given in Table 2. For the COX analysis, the total effect is given which consists of the aggregate main effects of $Privileged$ and the transition interaction ($y=0\rightarrow y=1)$, and likewise for $Contrast$. It can be seen that the GLM and MA1 methods provide similar estimates of experimental effects, and the AR1 and COX methods also have similar estimates. The LAG model incorporating the lag-1 $y$ predictor stands out against the other four models. In particular, the LAG coefficients for $Time$ and $Contrast$ are smaller, and the coefficient for $Privileged$ is higher. A ContrastTimeContrast*Time interaction is also evident for the lag model, and the $Time$ coefficient is strikingly dissimilar to those of the other models. The GLM, AR1, and MA1 models indicate a strong positive relationship between time and the likelihood of being on-target. The variable $Time$ is implicitly contained in the survival analysis model through run start and stop times. For this reason, the COX method did not produce intercepts or time interactions.
It is interesting that the AR1 and MA1 models resulted in moderately different estimates of experimental effects. In this regard, the MA1 model is most similar to the GLM model. Recall that the AR1, MA1, and COX approaches employ variance modeling alternatives to error correlation, while no adjustment for autocorrelation is present in the GLM model. Both models also indicate negligible interactions over time for the experimental variables.
7.2 Standard Errors
Standard errors for the fixed effects are given in Table 3. The models provide notably dissimilar SEs. The GLM provides the lowest, and the AR1/MA1, COX models provide the highest for the experimental valriables. The LAG model produces SEs about double of those for the GLM model, but even so, LAG SEs are half the size those for AR1, MA1, and COX. As noted earlier, the COX model has a lower effective sample size because the last run in each individual time series is dropped due to the lack of a transition indicator. Still, the SEs for the AR1, MA1, and COX models are very similar. This may imply that these three models implicitly weight transition nieghborhoods higher than other areas of the time series.
Table 3: Standard errors of fixed-effect estimates for five methods.
Term GLM AR1 MA1 LAG COX
Intercept 0.035 0.014 0.016 0.039
Ylag-1 0.012
Privileged 0.007 0.034 0.037 0.014 0.030
Contrast 0.004 0.021 0.023 0.009 0.018
Time 0.003 0.014 0.015 0.006
Priv*Time 0.007 0.033 0.035 0.015
Contrast*Time 0.005 0.020 0.022 0.009
  • Note. Robust SEs are reported for total effects for the COX model.
7.3 Random Effects and $\phi$
Random effects are given in Table 4 for the GLM and LAG models. For the AR1, MA1, and COX models, the cluster option was used in lieu of adding random effects. The estimated variance components obtained for the GLM and LAG methods were similar. For the AR1 and MA1 model, the estimated residual correlations are extremely high. However, this is not unexpected given the median run length of 35 (out of 112). This level of $\phi$ for AR1 and MA1 raises the concern that other coefficients may be destabilized. This is also true for the LAG model because a large pseudo $R^{2}$ is implied by the high magnitude of the $Ylag1$ coefficient.
Table 4: Random-effects variances and autoregressive parameters.
Model Subject variance Item variance $\phi$ $\alpha$
GLM 0.071 0.069
LAG 0.107 0.074
AR1 0.95 (0.948) 0.993
MA1 0.95 (0.963) 1.001
  • Note. Ridge estimates are reported for AR1 and MA1 with $\phi=.95$. Estimated values of $\phi$ are given in parentheses.
8 Discussion
The results of this study may shed some light on the issue of autocorrelation in eye-tracking data. Relative to the GLM and LAG models, AR1/MA1 and COX provide more conservative alternatives for estimating fixed effects. The results of the AR1 and COX models raise important and unresolved questions. First, the finding that these two approaches provide similar estimates of experimental effects may not be a coincidence. Recent work has shown that binary time–series models with autoregressive structure can behave very much like discrete-time two–state transition or survival models. For example, [campajola2020ising] demonstrated that a binary autoregressive process is mathematically equivalent to a Markov two–state model. Similarly, [gorgi2016survival] demonstrated how autoregressive recurrences can be expressed in terms of survival or transition probabilities. These relationships may help to explain why fitting an autoregressive model to a binary sequence of blocks can produce coefficients that resemble those obtained from modeling $0\rightarrow 1$ transitions.
Of course, the similarity of the AR1 and COX models does not imply they are correct. Rather it raises the question of how eye-tracking time series can be modeled. The AR1 and MA1 models are impractical for longer time series because the determinant of $R$ approaches zero. While a upper boundary can be set manually, this is purely an ad hoc solution. The survival model, or similar models, may provide a way to avoid near singularities due to serial correlation. Moreover, these models are not truly time series models but rather error-correlation models. Based on results from the current study, it is not clear that time series models are require or even suitable for eye-tracking data.
In hierarchical time-series models, it is well documented that autoregressive parameters and random effects are difficult to identify simultaneously. The coefficients $\phi$ of AR1/MA1 and $\beta_{0}$ of the LAG model can account for the same temporal structure as the subject and item random intercepts. When the available information is insufficient to distinguish their contributions, the model may distribute this dependence ambiguously between them, producing weak identification of $\phi$ and $\beta_{0}$, and unstable or inflated variance estimates for the random effects [diggle2002, galecki2013]. These difficulties arise particularly when individual time series are short [verbeke2000, shumway2017, jones1993]. For these reasons, methodological guidance commonly recommends avoiding the joint modeling of autocorrelation and random effects unless the study design provides strong and independent information to identify both components [pinheiro2000]. It is possible that a similar problem arises from modeling with a lag-1 dependent variable. For instance, in Table 2 the lag-1 coefficient seems to have redistributed $Time$ and $Contrast$ effects to the ContrastTimeContrast*Time interaction the subject variance component as shown in Table 4.
Data pre-processing is another important issue. Eye-tracking systems commonly compress data to avoid creating mammoth output files. Yet even with preliminary machine compression, files may be large. The RLE encoding method for additional compression is convenient for survival analysis and greatly reduces execution time. Pre-processing also includes data cleaning which involves tricky decisions of which experimental records to retain for analysis. This is especially the case in eye-tracking studies because momentary distractions can result in several consecutive seconds of invalid responses.
The results of this study collectively support the existence of experimental effects for the variable $Contrast$. For the AR1, MA1, and COX models, the effect for $Privileged$ is marginally significant. Also, the interaction terms for GLM and AR1/MA1 indicate the effects of $Privileged$ and $Contrast$ are uniform over time. Because estimates of the experimental effects in the current study varied in statistical strength, this demonstrates model assumptions can make substantive differences in both effect size and statistical significance.
The current study is based on a single data set, yet it does raise conceptual issues in the statistical modeling of eye-tracking data. Results are limited to a two-regime model for on-target and off-target, and a single data set. However, useful directions are suggested for future research on modeling issues in eye-tracking data. In particular, Markov transition models are suggested, but suitable software functions for the current data set would require capability for incomplete subject-by-item designs as well as clustering.
9 Acknowledgments
I used AI tools in the preparation of this paper. This included preparation of references in biblatex format, review of relevant psychological literature for understanding experimental variables, querying regarding modeling issues, suggesting R code for data file preparation, and obtaining latex code for tables and equations. I also generated three paragraphs of text which were heavily edited prior to finalizing the paper.